### v2

# LongAudio-RAG: Event‑Grounded Question Answering over Multi‑Hour Long Audio

Abstract
Long-duration audio is increasingly common in industrial and consumer settings, yet reviewing multi-hour recordings is impractical, motivating systems that answer natural‑language queries with precise temporal grounding and minimal hallucination. Existing audio–language models show promise, but long‑audio question answering remains difficult due to context‑length limits. We introduce LongAudio-RAG (LA-RAG), a hybrid framework that grounds Large Language Model (LLM) outputs in retrieved, timestamped acoustic event detections rather than raw audio. Multi‑hour streams are converted into structured event records stored in an SQL database, and at inference time the system resolves natural‑language time references, classifies intent, retrieves only the relevant events, and generates answers using this constrained evidence. To evaluate performance, we construct a synthetic long‑audio benchmark by concatenating recordings with preserved timestamps and generating template‑based question–answer pairs for detection, counting, and summarization tasks. Finally, we demonstrate the practicality of our approach by deploying it in a hybrid edge–cloud environment, where the audio grounding model runs on‑device on IoT‑class hardware while the LLM is hosted on a GPU‑backed server. This architecture enables low‑latency event extraction at the edge and high‑quality language reasoning in the cloud. Experiments show that structured, event‑level retrieval significantly improves accuracy compared to vanilla Retrieval-Augmented Generation (RAG) or text‑to‑SQL approaches.
LongAudio-RAG: Event‑Grounded Question Answering over Multi‑Hour Long Audio
Vakada Naveen, Kartik Hegde, Arvind Krishna Sridhar, Yinyi Guo, Erik Visser Qualcomm Technologies Inc., nvakada@qti.qualcomm.com, karthegd@qti.qualcomm.com, arvisrid@qti.qualcomm.com, yinyig@qti.qualcomm.com, evisser@qti.qualcomm.com
1 Introduction
Figure 1: Chat Example for LongAudio-RAG
Long-duration audio streams are now common not only in industrial settings such as machine monitoring and safety logging but also in homes through smart assistants, baby monitors, and security systems. These recordings span many hours, making manual review impractical and motivating systems that can answer natural-language questions about events and their timing. Unlike short-clip audio tasks, long audio question answering (QA) must handle time-bounded queries, aggregate counts, and narrative summaries, requiring precise temporal grounding and hallucination-free language generation.
Recent progress in audio-language modeling has enabled instruction-following and reasoning over audio using large-scale (audio, question, answer) training Chu et al. (2023); Huang et al. (2024); Gong et al. (2024); Ghosh et al. (2024); Goel et al. (2025). Despite these advances, long audio QA remains difficult in practice. First, raw multi-hour audio cannot be directly ingested by most models due to context-length and computation constraints, encouraging approaches that compress, segment, or selectively attend to relevant evidence. Second, natural-language time expressions are highly variable (12-hour vs. 24-hour time, shift references, relative phrases such as ‘before 5pm’), and errors in time interpretation can invalidate downstream reasoning. Third, open-ended generation over long logs is prone to hallucination unless responses are explicitly grounded in verifiable evidence.
Retrieval-augmented generation (RAG) reduces hallucinations by grounding model outputs in retrieved evidence Lewis et al. (2020); Gupta et al. (2024). Although widely used for knowledge-intensive QA and summarization, RAG performance is tightly coupled to retrieval quality, especially for ambiguous or underspecified queries Karpukhin et al. (2020); Gupta et al. (2024). Recent work has extended these ideas beyond text: audio-integrated RAG retrieves directly from audio–text representations rather than relying solely on transcripts Chen et al. (2025), while long-context video understanding similarly benefits from retrieval-based selection of relevant segments to manage large input spaces and improve faithfulness Luo et al. (2024); Tan et al. (2025).
This work presents a hybrid framework for long audio question answering that grounds LLM outputs in timestamped acoustic detections rather than generating responses directly from multi-hour audio, for chatbot application as shown in Figure 1. The system first converts long audio into a sequence of detected events with event name, timestamps, confidence scores and related attributes, stored in a SQL database for efficient retrieval. At query time, temporal expressions are resolved into concrete intervals, intent is classified (e.g., detection, counting, summary), and only the relevant events are passed to the LLM for answer generation as shown in Figure 2.
To enable controlled evaluation of time-bounded queries missing in existing short-audio datasets Wijngaard et al. (2025); Gemmeke et al. (2017) we construct a synthetic long audio benchmark by concatenating industrial and external recordings while preserving event timestamps. Template-based question–answer pairs are generated across detection, counting, and summary tasks with diverse temporal expressions, allowing systematic testing of long audio reasoning.
Contributions.
We present a comprehensive solution to the open problem of performing reliable question answering over multi‑hour audio recordings that can span across days.
  • (i)
    We introduce a hybrid grounding framework that anchors answers in timestamped acoustic events rather than raw audio, enabling precise temporal reasoning and reducing hallucination on hours-long content.
  • (ii)
    We proposed a complete, reproducible implementation stack, including event extraction, SQL-backed retrieval, temporal reference resolution, intent classification, and evidence-constrained generation.
  • (iii)
    We deploy the system in a hybrid edge-cloud configuration: audio grounding runs on IoT devices while the LLM is hosted on a GPU server, a split that reduces bandwidth usage, and scales language reasoning.
  • (iv)
    We design the system to be latency-aware by performing event detection and filtering at the edge, ensuring that only compact, relevant evidence is transmitted and enabling fast, stable responses even for multi-hour queries.
  • (v)
    We proposed a synthetic long audio benchmark and demonstrated improvements over text-to-SQL and RAG-based approaches.
To our knowledge, this is among the first event-level, edge–cloud systems for multi-hour audio question answering.
Figure 2: LongAudio-RAG (LA-RAG): Proposed method for long audio question answering.
2 Related Work
Research on retrieval‑augmented generation (RAG) highlights the value of grounding LLM outputs in external information to reduce hallucinations and improve factuality. Canonical RAG pipelines pair dense retrievers with generators to condition responses on retrieved passages Lewis et al. (2020), with surveys documenting challenges such as retrieval ambiguity, domain shift, and bias in retrieval corpora Gupta et al. (2024). Dense retrievers like DPR enable semantic matching beyond token overlap, but often struggle when queries are underspecified Karpukhin et al. (2020), motivating hybrid retrieval strategies and query refinement. Extending RAG beyond text has gained traction in audio and video domains, where raw modalities exceed model context windows. Audio‑focused approaches range from ASR‑dependent pipelines to end‑to‑end systems such as WavRAG, which retrieves from audio–text hybrid knowledge without requiring transcription Chen et al. (2025). Strong pretrained encoders such as wav2vec 2.0 remain crucial for robust audio retrieval and reasoning Baevski et al. (2020). In videos, RAG-like methods retrieve auxiliary OCR/ASR/object metadata or select question‑relevant temporal segments, as in Video‑RAG Luo et al. (2024) and RAG‑Adapter Tan et al. (2025). Efficient video‑to‑text conversion frameworks like ViTA Arefeen et al. (2024) and spatiotemporal architectures including I3D Carreira and Zisserman (2017) and TimeSformer Bertasius et al. (2021) further support long‑video retrieval and understanding.
Audio‑language modeling and audio question answering (AQA) focus on aligning acoustic events with natural‑language queries and often require fine-grained temporal reasoning. Early diagnostic datasets such as DAQA probe temporal and logical understanding through synthetic event sequences Fayek and Johnson (2020). Recent audio‑language models (ALMs) combine audio encoders with LLMs and rely on large‑scale QA datasets to enable open‑ended reasoning, with LTU trained on OpenAQA‑5M Gong et al. (2024) and GAMA introducing CompA‑R for complex temporal metadata reasoning Ghosh et al. (2024). Multi‑turn interaction datasets such as Audio Dialogues Goel et al. (2024) and multimodal instruction‑tuning resources like MULTIS Zhao et al. (2023) expand conversational and multimodal capabilities. Scaling multimodal systems to handle long-duration audio remains an open research challenge. Audio Flamingo-3 Goel et al. (2025); Yang et al. (2024)demonstrates notable progress through curriculum training and the use of curated long audio datasets, enabling support for audio inputs of up to 10 minutes , a limit that still falls short of the extended-duration requirements of our target use cases . In contrast, Qwen-3 Omni further advances long-context audio modeling by supporting audio inputs of up to 45 minutes, but still it doesn’t scale to the length we support Jin Xu and others (2025). Surveys cataloging audio‑language corpora note significant dataset overlap, bias, and language imbalance Wijngaard et al. (2025), with many resources grounded in large event-labeled corpora such as AudioSet Gemmeke et al. (2017). Finally, text‑to‑SQL systems Hong et al. (2025); Deng et al. (2021); Wang et al. (2025), can translate natural language queries into SQL commands that can be executed over traditional structured databases.
3 Methodology
Table 1: Evaluation results on the Simple-QA pairs ($n=800$) from the Home-IoT and Industrial-IoT datasets using various baseline models.
Data Audio Encoder Approach Detection (%) (n=300) Counting (%) (n=300) Summary (%) (n=200) Overall (%) Latency (s)
Home IoT AF3 RAG 49.87 53.73 23.60 44.75 5.77
AGM RAG 67.93 44.13 27.70 48.95 3.26
AGM Text2SQL 47.13 48.00 24.40 41.77 1.34
AGM LA-RAG (Ours) 90.67 76.93 56.10 76.88 0.56
Industrial IoT AF3 RAG 46.51 49.83 24.80 42.17 5.43
AGM RAG 66.00 47.47 29.30 49.88 3.66
AGM Text2SQL 48.07 44.53 24.70 40.90 1.08
AGM LA-RAG (Ours) 90.07 62.60 46.70 68.92 0.44
3.1 Automated QA generation
We generate evaluation QA pairs using a two-tiered approach that captures both controlled and realistic query variability for long‑audio understanding systems. Simple-QA pairs are created deterministically from ground‑truth event annotations produced by the Audio Grounding Model (AGM) module, covering detection, counting, and summary queries expressed in an unambiguous HH:MM:SS format. We created 3 types of queries - detection, counting, summary . To better reflect real‑world usage, we also produce complex-QA pairs that incorporate diverse temporal expression types including 24‑hour and 12‑hour formats,shift‑based references, before/after constraints, duration‑based phrasing, and other relative or segment-based intervals. More details about the QA pairs can be found in Appendix E.
3.2 Proposed approach for Long Audio Question Answering
First we process a long audio file by running our Audio Grounding Model (Section 3.2.1). using shifted windows to generate a text log of metadata ( Eg: audio event name , start time etc) as shown in Figure 2. This metadata is then inserted into a SQL database that can be queried. Then we retrieve the most relevant information from this database to answer the query. The detailed step by step process to generate the answering using the proposed method is given below:
3.2.1 Audio Grounding Model (AGM)
To preserve the flexibility required for on-device deployment across diverse acoustic environments and application scenarios, we develop an open-vocabulary Sound Event Detection (SED) model based on text to audio grounding. Instead of relying on a fixed label set, the model localizes sound events conditioned on free-form textual queries corresponding to the sounds of interest for a given use case. Our Audio Grounding Model (AGM) is trained following the phrase level WSTAG framework described in Xu et al. (2024), using the AudioCaps Kim et al. (2019) dataset from which we extract phrase-level supervision. AGM adopts the same architecture and hyperparameters as Xu et al. (2024): an audio encoder built from a CRNN with eight convolutional layers followed by a bidirectional GRU (BiGRU), and a text encoder composed of a single word embedding layer with mean pooling. Frame-level grounding scores are computed via the cosine similarity between audio frame embeddings and text tag embeddings, followed by a sigmoid activation. The output of this module is stored in JSON format which is then inserted into a SQLite database for question answering. The columns of the database are: audio event name, start time , end time , confidence of the generated class and other attributes like loudness.
3.2.2 Question Answering method
(1) Query Rephrasing: User queries are rewritten into clearer and more explicit forms by incorporating chat history, clarifying missing details, and stabilizing phrasing. This step reduces hallucinations, improves consistency, and strengthens downstream intent classification.
(2) Time Resolution: Natural-language time expressions are mapped to precise intervals through a rule-based extractor that supports explicit ranges, before/after constraints, duration-based phrasing, and configurable shift references. Non-temporal requests default to full-day intervals. When phrasing falls outside the rule set, an LLM-prompting based fallback method produces the start-end pair.
(3) Intent Classification: A hybrid classifier identifies whether the query concerns summarization, detection, counting, or anomaly analysis. Keyword-based detection provides fast matches, while embedding based similarity captures indirect or paraphrased intents, enabling accurate routing across diverse query styles.
(4) SQL based retrieval and LLM based Response generation: After rephrasing, time resolution, and intent detection, the system retrieves relevant events from the database of audio events metadata. Broad queries use the full event set, whereas specific ones apply an embedding similarity based Tan et al. (2019) Top-k filter to use the most relevant audio events . An intent-specific prompt template (summary or detection/counting) then conditions the LLM on filtered, timestamped events to generate grounded and faithful responses. Anomaly queries follow a separate specialized workflow (Appendix I). More details on all these steps along with the prompts used can be found in Appendix F,G and M.
4 Results
We evaluate our system across multiple large language models (LLMs), data-processing pipelines, and query types to assess its effectiveness in real-world Industrial IoT audio analytics. The evaluation focuses on three downstream tasks derived from machine logs: event detection, event counting estimation, and natural language summarization.
4.1 Dataset for evaluation
We built a controlled synthetic benchmark using two 24-hour audio recordings, one for home IoT (HIoT) and one for industrial IoT (IIoT) by concatenating short, labeled events to preserve timestamps and vary density and noise (event classes listed in Appendix B). To match evaluation scope, we restrict acoustic classes to these predefined lists and pass the same closed set to downstream prompts, enabling grounded reasoning over multi-hour timelines without drift. Background audio uses industrial white-noise machine recordings for IIoT and low-hum ambience for HIoT, sourced from Freesound Fonseca et al. (2021). The final test vectors have a global SNR of 6 dB and are normalized to the Apple loudness target of –16 LUFS.
Table 2: Comparison Across Model Scales on Industrial-IoT dataset
Model Name # Active Params Detection (%) counting (%) Summary (%) Overall (%)
Phi-3-mediumArah A (2024) 14B 93.13 63.47 48.60 70.88
Minitron Muralidharan et al. (2024) 8B 90.67 63.67 42.70 68.55
Qwen3-8B Yang et al. (2025) 8B 86.20 43.53 36.90 57.88
Llama-3.1Grattafiori et al. (2024) 8B 60.60 38.80 26.30 43.85
Qwen2.5 Yang et al. (2024) 7B 90.73 64.60 38.50 67.88
Phi-3.5-MoE Arah A (2024) 6.6B (42B) 92.07 64.93 48.90 71.10
Phi-4-mini-instruct Abouelenin et al. (2025) 3.8B 92.60 61.60 43.80 68.77
Llama-3.2 Grattafiori et al. (2024) 3B 51.67 38.27 28.40 40.83
Qwen2.5 Yang et al. (2024) 0.5B 70.80 41.67 38.20 51.74
Table 3: Time resolution module evaluation results
Difficulty Regex-only (%) LLM-only (%) Combined (%)
Easy 93.33 86.67 100.00
Medium 55.00 65.00 85.00
Hard 20.00 20.00 30.00
Overall 60.00 62.22 77.78
4.2 Evaluation metrics
We evaluated the system’s performance across three question categories using GPT-4o as an automated judge with category-specific rubrics on a 1-5 scale. For detection questions (yes/no queries), the evaluator primarily assessed answer correctness while treating explanations as optional enhancements, with correct answers scoring 4-5 and incorrect answers capped at 2 regardless of explanation quality. Counting based questions were evaluated based on numeric accuracy, where exact matches scored 4-5, off-by-one or within-10% errors scored 3, and order-of-magnitude errors scored 1. Summary questions received the most nuanced evaluation, considering content coverage, factual accuracy, completeness, and coherence, where scores of 5 required all key points with accurate temporal/causal relationships, while scores of 3-4 indicated partial coverage with minor omissions or inaccuracies.
4.3 Baseline methods for comparison
RAG Baseline
The retrieval baseline encodes each audio event as a concise text record containing its tag, timestamp, duration, and loudness. At query time, semantically relevant events are retrieved, and the model answers using only these snippets, allowing us to evaluate QA performance without explicit structured reasoning. We compare two variants by swapping our Audio Grounding Model with Audio Flamingo 3 (AF3) as the audio encoder. In the AF3-based RAG pipeline, we prompt the model to generate event names, timestamps, and scene descriptions, but it fails to produce reliably structured outputs. Instead, audio is processed in 4-minute chunks and stored in a vector database, with $k=5$ used for all RAG experiments. We used sentence-transformerReimers and Gurevych (2019) based all-MiniLM-L6-v2 as the embedding model for all RAG experiments.
Text-to-SQL Baseline
The text-to-SQL baseline treats long‑audio question answering as a structured querying task. An LLM is prompted to convert the user query into an SQL query that can be executed over the database. All audio events are stored in a relational table, and natural‑language questions are translated into SQL that filters, aggregates, or ranks events based on their tags and temporal attributes. The final answer is generated by interpreting the resulting table, allowing this baseline to test how well symbolic querying captures fine‑grained, time‑specific information in long‑duration audio.
4.4 Audio Grounding Model evaluation
We evaluate the standalone AGM on long form audio using the Sound Event Detection (SED) task using DESED datasetSerizel et al. (2020) from the AudioMarathon benchmarkHe et al. (2025), a recent suite designed to assess long context audio understanding. For each multiple choice question (MCQ), we pass the four candidate sound event tags to AGM and compute a score for each tag by averaging the model’s frame level similarity scores across the entire audio clip. The predicted answer corresponds to the tag with the highest aggregated score. Under this simple inference setup, AGM achieves an F1 score of 74.8, ranking just below the top performing Qwen2.5 Omni 7B model, which attains 78.4 on the leaderboard ( Appendix D ). To obtain temporal event boundaries from AGM output, we apply a global decision threshold of 0.8 to the framewise scores and refine the resulting predictions using a 0.3 second median filter, which we hold constant across all experiments.
4.5 Comparison of Querying Approaches
We evaluate four approaches for querying Industrial IoT audio: (1) our method: AGM + longAudio-RAG, (2) AGM + RAG, (3) AGM + Text2SQL over event tables, and (4) AF3 + RAG. As shown in Table 1, our approach consistently outperforms the others across detection, counting, and summary tasks. By using structured AGM logs, explicit temporal resolution, and intent-aware processing, the system delivers more accurate detections and more reliable event-level reasoning than retrieval- or SQL-based pipelines.
Standard RAG struggles on detection and counting because retrieved text lacks precise temporal structure, and LLMs are unreliable at boundary-aware operations. Text2SQL also often fails, producing invalid queries or misinterpreting time intervals, resulting in incomplete or incorrect retrieval.
To isolate the impact of high-quality event extraction, we compare against an AF3-based RAG pipeline. AF3 + RAG performs substantially worse, reinforcing that (i) well-structured AGM event logs are essential for downstream temporal reasoning, and (ii) retrieval grounded in event-level representations is more effective than operating on AF3 features or transcript-level content. We additionally conducted a human evaluation study using a curated set of 113 questions, divided into three categories: 67 detection, 35 counting, and 11 summarization questions. A human evaluator rated system responses on a 1–5 scale, on which our approach with Phi-3.5-MoE LLM, achieved an overall average score of 4.28. Further details on the human and GPT‑based evaluation setup and analysis are provided in Appendix H.
4.6 LLM Comparison Across Model Scales
Table 2 compares LLMs from 0.5B to 14B active parameters, showing that larger models consistently perform better, especially on detection and summary tasks that require precision and contextual reasoning. The strongest results come from Phi-3.5-MoE (71.10%; 42B, 6.6B active) and Phi-3-medium (70.88%; 14B), these models show strong reasoning and reliable interpretation of the generated SQL tables. Models in the 7-8B range, such as Minitron-8B and Qwen2.5-7B, also perform competitively with scores near 68%. In contrast, smaller models ($\leq$3B), including Llama-3.2-3B and Qwen2.5-0.5B, show clear degradation across tasks, indicating that adequate model capacity is essential for structured-context understanding.
4.7 Time resolution module analysis
To evaluate our time‑resolution methods, we created 45 QA pairs spanning difficulty (easy, medium, hard) and semantic type (details in Appendix F). Easy questions use explicit, unambiguous expressions (e.g., “between 2pm and 4pm”), medium questions require interpreting relative phrases and simple arithmetic (e.g., “first 2 hours of day shift”), and hard questions involve ambiguous or colloquial references, boundary cases, or multi‑step constraints (e.g., “around lunchtime”). As shown in Table 3, the combined regex+LLM approach achieves the highest overall accuracy (77.78%), including perfect accuracy on easy cases and strong improvements on medium and hard queries.
4.8 Deployment Considerations
The system has been deployed in real-world environments with all components running in production. The Audio Grounding Model (AGM) operates on the Qualcomm IQ-9075 platform, which delivers up to 100 TOPS of on-device AI performance, enabling low-latency audio processing while keeping raw audio local for privacy. Its octa-core Kryo Gen6 CPU, Hexagon Tensor Processors, and industrial operating range support continuous edge inference. The LLM and embedding services run on NVIDIA A100 GPU servers, providing the throughput required for large-scale embedding generation and model inference. Other details of the software implementation can be found in Appendix K.
4.9 Inference Latency
Table 1 reports the end-to-end average latency per query (seconds) for all Phi-4-mini-instruct configurations across Home IoT (HIoT) and Industrial IoT (IIoT) scenarios. Our proposed method is the fastest, Text-to-SQL incurs modest overhead, RAG is slower, and AF3 prompt-based runs remain the most expensive.
5 Conclusion
We presented an end-to-end framework for multi-hour audio understanding that integrates a lightweight AGM, structured event logs, and LLM-based reasoning. Our approach outperforms standard RAG and Text2SQL pipelines by leveraging temporally aligned AGM events for accurate detection, counting, and summarization. Experiments across diverse datasets show that mid-sized instruction-tuned and MoE models, when paired with high-quality event representations, deliver strong performance without relying on very large LLMs. Overall, the results demonstrate that structured event extraction combined with efficient retrieval forms a reliable foundation for machine-centric audio analytics. Future work includes expanding to additional sensing modalities and introducing agentic capabilities that leverage tools such as text-to-SQL and code generation. We also aim for fully end-to-end on-device deployment, running both the AGM and compact LLMs on edge hardware for low-latency inference.
Appendix A Ethical Considerations
This work involves processing long-duration audio, which may contain sensitive personal or environmental information. To mitigate privacy risks, all raw audio is processed locally on edge devices, and only structured event metadata is transmitted to the server, reducing the likelihood of exposing identifiable content. Our system does not attempt to infer personal attributes, emotions, or identities, and all event logs used for evaluation are synthetic or derived from publicly available datasets to avoid handling real user data. Finally, we ensure that model outputs are grounded in retrieved evidence to minimize hallucinations and reduce the risk of generating misleading or harmful responses.
Appendix B Audio classes for Home-IoT and Industrial IoT datasets
The home set covers alarms, sirens, door_bell, door_knock, glass_breaking, car_crash, door_close-open, baby_cry, gun_shot, cat, car_honk, snoring, dog_bark; the industrial set includes tools clanking, hand saw, hand file, workers talking, footsteps, arc welder, diesel forklift, power hand drill, stamping machine, walkie talkie, warning buzzer, factory whistle
Appendix C Custom sound enrollment and detection
Our on-device long audio QA framework also supports custom sound enrollment, allowing users to upload example clips of sound events of interest and query for them in continuous audio. To enable this capability, we train a prototypical network on FSD50K Wu et al. (2023) using CLAP Fonseca et al. (2021) audio embeddings, formulated as an N way M shot classification task where the model learns to discriminate among N classes given M support samples. To preserve the CLAP audio encoder, we freeze all of its parameters and learn a single 512 dimensional linear projection layer for custom sound adaptation. Training uses negative log-likelihood loss with softmax scoring, and classification is based on cosine similarity normalized by a temperature scaled LogSoftmax. At inference time, we apply a 5 second sliding window with 4 second overlap, yielding a per second score for each enrolled sound; a global threshold of 0.5 is then applied to these framewise scores to determine temporal event boundaries.
Appendix D AGM evaluation
Table 4 reports Sound Event Detection (SED) performance on the AudioMarathon benchmark , comparing leading long‑form audio understanding models. Our AGM model, evaluated on the DESED dataset, achieves a strong F1 score of 74.8, ranking just below the top Qwen2.5‑Omni‑7B system.
Table 4: Sound Event Detection (SED) Results on AudioMarathon Benchmark. The entry marks with ’*’ is our result. The remaining entries are cited from He et al. (2025)
. Rank Model SED (F1) 1 Qwen2.5-Omni-7B 78.4 2 AGM (Ours)* 74.8 3 Voxtral-Mini-3B-2507 71.0 4 Qwen2.5-Omni-3B 70.2 5 Gemma-3n-E2B-it 50.2 6 Gemma-3n-E4B-it 50.2 7 Audio-Flamingo-3 59.5 8 Phi-4-Multimodal 55.1 9 Aero-1-Audio 55.0 10 Baichuan-Omni-1.5 45.7 11 Audio-Flamingo-2 27.1
Appendix E Automatic QA pair generation
E.1 Simple-QA pair generation
The development of robust evaluation frameworks for long audio based chatbot systems necessitates the creation of comprehensive question-answer (QA) datasets that can systematically assess model performance across diverse query types and temporal expressions. In this work, we present a simple yet effective QA generation methodology that employs a deterministic approach to create evaluation benchmarks for long-form audio understanding systems. Our framework generates three primary categories of queries: detection queries (binary presence/absence verification), counting queries (occurrence counting), and summary queries (temporal and statistical aggregation). To ensure temporal consistency and reproducibility, we adopt a standardized HH:MM:SS time format for all temporal expressions, eliminating ambiguities inherent in natural language time references such as "morning shift" or "12-hour format" variations. The generation process leverages ground-truth audio event annotations obtained using Audio Grounding Model (AGM) to construct queries with verifiable answers, incorporating both positive samples (events present in the audio) and negative samples (absent events and unrelated event categories) to evaluate false positive rates. Furthermore, we introduce synonym-based query variations to assess the semantic understanding capabilities of the target system, replacing canonical event labels with contextually equivalent terms. This simple generation approach produces balanced datasets with controlled distributions across query categories, enabling systematic evaluation of audio QA systems while maintaining interpretability and reproducibility, critical requirements for benchmarking in the audio understanding domain.
E.2 Complex-QA pair generation
While simple-QA generation provides a foundational evaluation framework, real-world audio understanding systems must handle the inherent complexity and variability of natural language temporal expressions. To address this challenge, we propose a complex-QA generation methodology that systematically incorporates diverse temporal reference formats to evaluate the robustness of audio event detection systems under realistic query conditions. Our approach extends beyond standardized HH:MM:SS formats to encompass seven distinct temporal expression types: (1) standard 24-hour format (e.g., "between 08:00 and 16:00"), (2) 12-hour format with meridiem indicators (e.g., "from 2:30 pm to 4:00 pm"), (3) shift-based references aligned with industrial or operational contexts (e.g., "during the morning shift"), (4) before/after temporal boundaries (e.g., "after 17:30"), (5) duration-based expressions relative to shifts or day boundaries (e.g., "first 2 hours of day shift"), (6) half-based temporal divisions (e.g., "second half of afternoon shift"), and (7) relative duration expressions (e.g., "between 1 hour and 2 hours"). By systematically varying temporal expression complexity, our QA generation facilitates the identification of model weaknesses in time parsing, semantic interpretation, and cross-format generalization, critical capabilities for deploying audio understanding systems in production environments where users employ diverse and unpredictable temporal references.
E.3 Implementation Details and Dataset Characteristics
Multi-Domain Support: The framework supports two configurable IoT domains: Industrial IoT (12 event classes: arc welder, stamping machine, diesel forklift, etc.) and Home IoT (13 event classes: baby cry, doorbell, glass breaking, etc.), each with domain-specific event taxonomies and synonym mappings to ensure contextual relevance.
Dataset Scale: The generation process produces 100 QA pairs per section across five generation phases: (1) detection & frequency with original event labels, (2) detection & frequency with synonym variations, (3) detection & frequency with unrelated events (negative samples), (4) specific event summaries, and (5) generic summaries. Since detection and frequency queries are generated in pairs for the first three sections, this yields approximately 800 total QA pairs (600 detection/frequency + 200 summary) with balanced positive/negative distributions and a minimum 10-second temporal window.
Temporal Expression Distribution: In complex-QA mode, temporal expressions are sampled using a probability-based strategy: 5% full-day queries, 15% shift-based (when enabled), 10% before/after boundaries, 10% duration-based, 5% half-based, 20% 12-hour format, and the remainder as standard 24-hour format, ensuring comprehensive coverage across temporal expression types.
Shift Definitions: For industrial contexts, standardized shifts are defined as: Morning/Day Shift (08:00–16:00), Afternoon/Evening Shift (16:00–24:00), and Night Shift (00:00–08:00). Shift-based queries can be toggled via a configuration flag, enabling evaluation with or without shift-specific temporal reasoning.
Ground Truth Structure: Each QA pair includes structured JSON metadata containing query category/subcategory, temporal bounds (HH:MM:SS), event tags, time expression type, and relevant statistics (detection status, occurrence counts, first/last event timestamps), enabling automated evaluation and error analysis.
Validation Mechanisms: The framework implements safeguards including configuration validation with automatic fallbacks, maximum attempt limits (10× target count) to prevent infinite loops, minimum duration enforcement (10 seconds), and automatic mode compatibility checks.
Appendix F Time resolution
The detailed category-level results in Appendix Table 5 highlight the specific categories that influence model performance. Categories with well-structured or explicitly defined temporal expressions such as explicit time ranges, shift-based cues, and full-day implicit references show consistently strong performance, particularly when using the combined approach. In contrast, categories involving greater variability or ambiguity, including relative durations, typos and variations, and edge cases, yield lower accuracy across all methods, reflecting their inherent complexity. This breakdown provides granular insight into error patterns and complements the difficulty-based analysis presented in the main paper. Detailed flowchart and prompt are shown in Figure 3 and 4.
Table 5: Category-wise accuracy results for the time resolution module.
Category Regex-only LLM-only Combined
explicit_time_ranges 100.00 100.00 100.00
shift_based 100.00 100.00 100.00
relative_durations 60.00 60.00 100.00
before_after 100.00 50.00 100.00
half_periods 100.00 50.00 100.00
full_day_implicit 100.00 100.00 100.00
typos_and_variations 27.27 72.73 72.73
edge_cases 20.00 20.00 30.00
Figure 3: Time resolution module
Figure 4: Prompt used in time resolution module
Appendix G Query rephrasing module
The detailed flowchart and the prompt used for query rephrasing module are shown in Figure 5 and 6.
Figure 5: Query Rephrasing Module
Figure 6: Prompt used in Query Rephrasing Module
Appendix H Dialogue evaluation using humans and GPT-4o as judge
We evaluated a dialogue system on 113 single‑turn QA pairs across three task categories detection, counting, and summarization using both a human evaluator and a GPT-4o based evaluator Achiam et al. (2023). For every response, each judge assigned 1–5 scores for language quality, correctness, reasoning coherence, and user experience; we averaged these four scores to obtain a per‑item composite and then averaged those composites within each category to produce the final category scores. Human evaluation yielded category means of 4.36, 4.25, and 3.95 (detection, counting, summarization), while GPT-4o based evaluation yielded 4.46, 3.67, and 3.47 for the same categories. The Pearson correlation between human and GPT per‑item composites across all 114 examples was 0.57, indicating moderate alignment. Despite comparatively lower outcomes in counting and summarization, performance is likely to improve with model scaling or targeted fine‑tuning within the size–accuracy budget. Overall, the system delivers strong conversational performance and coherent execution across categories
Appendix I Anomaly Detection Extension
Our framework naturally supports anomaly detection by treating anomalies as structured event-level signals that can be retrieved, analyzed, and explained using the same time-resolution and intent-routing components. When users ask anomaly-related questions, the system first classifies the query subtype (e.g., loudness anomaly, pitch anomaly, or abnormal start-time), retrieves the relevant events, and computes anomalies using a plug-in detector. In our implementation, we evaluate two forms of attribute-based anomalies such as loudness deviations from the normal range and temporal anomalies, where event start times fall outside expected intervals derived from historical clusters. The system produces a structured anomaly table and, when needed, incorporates guidance from user-provided instruction manuals before prompting the LLM to generate a grounded explanation. This allows the method to flexibly support any anomaly detection backend while maintaining faithful, event-grounded audio question answer capabilities.
We have implemented anomaly detection based on the loudness attribute and also irregular sound patterns. Figures 10 and 8 show an example with the flowchart analysis pipeline that can be integrated into our proposed system. Figure 9 shows how we can integrate user manuals and other custom logs into our system for more personalized responses to target the specific use-case.
Figure 7: Loudness anomaly detection
Figure 8: Frequency based anomaly detection
Figure 9: Anomaly detection processing for long audio QA
Appendix J Limitations
Our system relies on the accuracy of the AGM module, so missed or noisy detections can propagate to downstream reasoning. The time-resolution component may also struggle with rare or highly ambiguous temporal expressions. Although our synthetic benchmark enables controlled evaluation, it does not fully capture the diversity of real-world acoustic environments.
Appendix K Software Implementation
The system is implemented as a modular, service-oriented stack comprising an edge AGM inference service, a backend indexing and QA service, and a lightweight Streamlit-based frontend. The AGM service runs a compact PyTorch model via FastAPI, producing JSON event logs with temporal metadata and supporting custom sound enrollment by allowing users to register new classes from example recordings. Its low computational footprint enables deployment on resource-constrained edge devices, reducing the need to transmit raw audio. The backend ingests these logs, stores them in a database, and builds retrieval structures for downstream reasoning; it initializes an embedding model for semantic indexing and an LLM for answer generation, orchestrated through LlamaIndex and served efficiently using vLLM on GPU machines. This backend also maintains conversational state for multi-turn queries. The frontend provides a simple interface for uploading audio, enrolling new sound classes, and issuing natural-language queries, communicating with the backend and inference service through RESTful APIs to allow each component to scale independently.
Figure 10: Software Implementation Architecture
Appendix L User Interface for LongAudio-RAG
We have implemented a web application front-end which allows the users to start recording the audio or upload an existing audio file and then chat over its contents. The UI also provides the option to enroll custom audio sounds that will be supported by the system. The screenshot of the UI is shown in Figure 11
Figure 11: LongAudio-RAG user interface
Appendix M Prompts used for intent based response generation
The prompts used for the response generation are shown in Figures 12 and 13.
Figure 12: Prompt for Summary category
Figure 13: Prompt for detection/counting based questions